OpenStudy (anonymous):

Help please What is the equation of the line perpendicular to –x + y = 7 and passing through (-1, -1)?

OpenStudy (photon336):

do you know how to write this in the form? \[y = m(x)+b\]

OpenStudy (photon336):

@McFly_47

OpenStudy (anonymous):

@Photon336 Is that the slope formula?

OpenStudy (photon336):

so yeah this is the formula for an equation of a line y = m(x)+b slope intercept form. since we've got -x+y = 7 we need to first get y by itself how do we do this?

OpenStudy (anonymous):

@Photon336 to be quite honest, I don´t know, can you help me?

OpenStudy (anonymous):

@Photon336 x – y = 2 is this the answer?

OpenStudy (photon336):

here is what we do we get y by itself and we do this by the following. see that -x we add it to both sides, doing the opposite operation. -x+y = 7 +x +x then we get y = x+7

OpenStudy (photon336):

\[y_{1} = x+7 \] now to find the line perpendicular to this, we need to take the -reciprocal of the slope. We have 1 so it would be -1 so far y = -x+? okay, so far so good ?

OpenStudy (anonymous):

so far so good sir

OpenStudy (photon336):

okay great. when the line must also pass through (-1,-1) have an idea of how to do this?

OpenStudy (anonymous):

To be honest, no. I´m only 14 an they put me in this class and it is hard @Photon336

OpenStudy (anonymous):

Can you teach me?

OpenStudy (photon336):

okay no problem so (-1,-1) means our x coordinate is -1 and our y coordinate is (-1)

OpenStudy (photon336):

so for the equation to pass through this point we must put the point in. \[y = m(x)+b \] \[y = -x+b\] \[(-1) = -(-1)+b\] we don't know what our b value is so we need to find it.

OpenStudy (anonymous):

umm so b would count as 1 right?

OpenStudy (anonymous):

and that I think would equal -2?

OpenStudy (photon336):

so let's check that

OpenStudy (photon336):

if b is -2 then our equation becomes this \[y = -x-2 \] now when we plug in -1 we should get -1 also -(-1)-2 1-2 = -1 so this works do you see this?

OpenStudy (anonymous):

I see how that works, but, would that be the answer? I have 5 options -x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0

OpenStudy (anonymous):

@Photon336

OpenStudy (photon336):

so look at the second one -x-y = 2 can you solve for y, or in other words get y by itself?

OpenStudy (anonymous):

Adding y to both sides and then dividing by 2?

OpenStudy (anonymous):

@Photon336

OpenStudy (photon336):

what would be our answer ?

OpenStudy (anonymous):

x=−y−2 ?

OpenStudy (anonymous):

is that right? @Photon336

OpenStudy (photon336):

watch what I do

OpenStudy (photon336):

-x-y = 2 we get y by itself. see the -x we add that to both sides and get -y = 2+x or -y = x+2

OpenStudy (photon336):

now after we do this, we need to get that - sign away from that y so it's actually this \[-1*y = x+2 \] so we divide everything by -1 \[y = -x-2\] look familiar?

OpenStudy (anonymous):

Yeah its like m answer but switching the x and y

OpenStudy (photon336):

yep so that's the answer

OpenStudy (photon336):

-x-y = 2

OpenStudy (anonymous):

@Photon336 wow, thank you so much for your time, and this looks like a simple problem but it sure is hard

OpenStudy (photon336):

I'll show you how to work with the formula it's pretty simple.

OpenStudy (anonymous):

@Photon336 I am so sorry my Internet went out, I would appreciate that

OpenStudy (photon336):

yeah, no problem

OpenStudy (anonymous):

@Photon336 Just one question if its not too much to ask, how is the formula called?